A boy throws a ball from shoulder height level at an initial velocity of 30 m/s. Spending 4.8s in air, the ball is caught by another boy as the same shoulder height level. What is the angle of projection? a) 37 b) 30 c) 53 d) 60

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Nikita
 
As we know
t = 2* u * sin@/g
 4.8= 2* 30 * sin@/10
4.8 = 6 sin@
48/60 = sin@
Sin@ = 4/5
Hence @ = 53°
 
Regards
Arun (askIITians forum expert)